# Pre-training Large Memory Language Models with Internal and External Knowledge

### v1

@@ Acknowledgments
SZ is supported by the Defense Advanced Research Projects Agency (DARPA) under Grant No. D24AP00259-00. CKB is supported by the National Science Foundation (NSF) through the NSF Research Traineeship (NRT) program under Grant No. 2345579. JPZ is supported by a grant from the Natural Sciences and Engineering Research Council of Canada (NSERC) (567916). This work is partially supported by LinkedIn-Cornell Partnership, Open Philantropy, the National Science Foundation NSF (IIS-2107161, and IIS-1724282, HDR-2118310), the Cornell Center for Materials Research with funding from the NSF MRSEC program (DMR-1719875), DARPA, arXiv, and the New York Presbyterian Hospital.
@@ Appendix C Detailed Results
No fine-tuning is performed; models are evaluated with retrieval at inference time only. In practice, we observe that the BM25 retriever consistently retrieves highly relevant supporting documents, ensuring a fair and competitive RAG baseline. We evaluate RAG with OpenAI/GPT2-355M, enabling a direct comparison against our LmLm model of similar scale.
@@ D.4 Qualitative Results
LLaMA2-176M-LmLm Ko Itakura is the <|db_entity|> Ko Itakura<|db_relationship|> Position<|db_return|> center-back, defensive midfielder<|db_end|> center-back and defensive midfielder for <|db_entity|> Ko Itakura<|db_relationship|> Current Team<|db_return|> Japan national team<|db_end|> Japan’s senior squad in the <|db_entity|> Ko Itakura<|db_relationship|> League Participation<|db_return|> 2022 FIFA World Cup<|db_end|> 2022 FIFA World Cup. He was born on <|db_entity|> Ko Itakura<|db_relationship|> Birth Date<|db_return|> January 27, 1997<|db_end|> January 27, 1997. Career. On <|db_entity|> Ko Itakura<|db_relationship|> Joined Club<|db_return|> Kawasaki Frontale<|db_end|> July 30, 2015, it was announced that Itakura joined JFL club Kawasaki Frontale from their youth system. In <|db_entity|> Ko Itakura<|db_relationship|> Moved To<|db_return|> Groningen<|db_end|> August 2016 he moved to Groningen. On <|db_entity|> Ko Itakura<|db_relationship|> Debut Date<|db_return|> June 17, 2019<|db_end|> June 17, 2019, Itakura debuted against <|db_entity|> Ko Itakura<|db_relationship|> Debut Opponent<|db_return|> Uruguay<|db_end|> Uruguay in an international friendly match. After playing only one game for Groningen during his debut season, Itakura left the club at the end of the year.
@@
LLaMA2-382M-LmLm Ko Itakura is the <|db_entity|> Ko Itakura<|db_relationship|> Position<|db_return|> center-back, defensive midfielder<|db_end|> center-back and defensive midfielder for <|db_entity|> Ko Itakura<|db_relationship|> Team<|db_return|> Japan national team<|db_end|> Japan’s national team. He was born in <|db_entity|> Ko Itakura<|db_relationship|> Birthplace<|db_return|> Yokohama<|db_end|> Yokohama on <|db_entity|> Ko Itakura<|db_relationship|> Birth Date<|db_return|> January 27, 1997<|db_end|> 27th January 1997. His father is from <|db_entity|> Ko Itakura<|db_relationship|> Father’s Origin<|db_return|> Japanese<|db_end|> Japanese and his mother is from <|db_entity|> Ko Itakura<|db_relationship|> Mother’s Origin<|db_return|> Japanese<|db_end|> Japanese. When he was young, he played baseball but after watching an exhibition match against a professional baseball club, he decided to become a footballer. In <|db_entity|> Ko Itakura<|db_relationship|> Joined Club Year<|db_return|> Kawasaki Frontale<|db_end|> 2013, he joined J1 League side Kawasaki Frontale. However, he could not play many matches behind Shusaku Nishikawa until <|db_entity|> Ko Itakura<|db_relationship|> First Match as Starter<|db_return|> Uruguay<|db_end|> September when he debuted at right back against Uruguay. After that, he became a regular player under manager <|db_entity|> Ko Itakura<|db_relationship|> Manager Under Whom Became Regular Player<|db_return|> Japan national team<|db_end|> Shinji Ono. On <|db_entity|> Ko Itakura<|db_relationship|> Debut Date<|db_return|> June 17, 2019<|db_end|> 17 June 2019, he debuted with Japan national team against Chile during the <|db_entity|> Ko Itakura<|db_relationship|> Competition Debut<|db_return|> Uruguay<|db_end|> 2019 Copa América. Career statistics. “Updated to end of 2018 season”. National team career. In August 2016, Itakura was elected to the <|db_entity|> Ko Itakura<|db_relationship|> U-23 Selection<|db_return|> Japan U-20 national team<|db_end|> Japan U-20 national team for the <|db_entity|> Japan U-20 national team<|db_relationship|> Tournament Participation<|db_return|> 2017 U-20 World Cup<|db_end|> 2017 U-20 World Cup. At this tournament, he played all 4 matches as left back of three back defense. In May 2019, he was selected for the <|db_entity|> Ko Itakura<|db_relationship|> Senior Squad Selection<|db_return|>

### v2

@@ Acknowledgments
SZ is supported by the Defense Advanced Research Projects Agency (DARPA) under Grant No. D24AP00259-00. CKB is supported by the National Science Foundation (NSF) through the NSF Research Traineeship (NRT) program under Grant No. 2345579. JPZ is supported by a grant from the Natural Sciences and Engineering Research Council of Canada (NSERC) (567916). This work is partially supported by LinkedIn-Cornell Partnership; Open Philantropy; the National Science Foundation NSF under awards OAC-2311521, IIS-2107161, IIS-1724282, and HDR-2118310; NASA under award No. 20-OSTFL20-0053; the Cornell Center for Materials Research with funding from the NSF MRSEC program (DMR-1719875); DARPA; arXiv; and the New York Presbyterian Hospital. Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation or of NASA.
@@ Appendix C Detailed Results
As an additional note on Table 2, RAG performs reasonably well on FactScore, and we expect further improvements with larger, instruction-tuned models, as smaller models may struggle to effectively use retrieved context. RAG also achieves high scores on T-REx, though this may reflect the benchmark’s overlap with Wikipedia, where many completions are retrieved verbatim. We include RAG results to provide a broader empirical context.
While both RAG and LmLm access external knowledge, they differ fundamentally. RAG retrieves unstructured text from Wikipedia and relies on the model to extract relevant content. In contrast, LmLm learns to issue explicit, structured lookups only when needed, interleaving retrieval with generation at the entity level. This enables more precise and easily verifiable access to factual knowledge. The two approaches are complementary—RAG could be applied on top of LmLm for potential additional gains.
@@ D.4 Qualitative Results
LLaMA2-176M-LmLm Ko Itakura is the <|db_start|> Ko Itakura<|sep|> Position<|db_retrieve|> center-back, defensive midfielder<|db_end|> center-back and defensive midfielder for <|db_start|> Ko Itakura<|sep|> Current Team<|db_retrieve|> Japan national team<|db_end|> Japan’s senior squad in the <|db_start|> Ko Itakura<|sep|> League Participation<|db_retrieve|> 2022 FIFA World Cup<|db_end|> 2022 FIFA World Cup. He was born on <|db_start|> Ko Itakura<|sep|> Birth Date<|db_retrieve|> January 27, 1997<|db_end|> January 27, 1997. Career. On <|db_start|> Ko Itakura<|sep|> Joined Club<|db_retrieve|> Kawasaki Frontale<|db_end|> July 30, 2015, it was announced that Itakura joined JFL club Kawasaki Frontale from their youth system. In <|db_start|> Ko Itakura<|sep|> Moved To<|db_retrieve|> Groningen<|db_end|> August 2016 he moved to Groningen. On <|db_start|> Ko Itakura<|sep|> Debut Date<|db_retrieve|> June 17, 2019<|db_end|> June 17, 2019, Itakura debuted against <|db_start|> Ko Itakura<|sep|> Debut Opponent<|db_retrieve|> Uruguay<|db_end|> Uruguay in an international friendly match. After playing only one game for Groningen during his debut season, Itakura left the club at the end of the year.
@@
LLaMA2-382M-LmLm Ko Itakura is the <|db_start|> Ko Itakura<|sep|> Position<|db_retrieve|> center-back, defensive midfielder<|db_end|> center-back and defensive midfielder for <|db_start|> Ko Itakura<|sep|> Team<|db_retrieve|> Japan national team<|db_end|> Japan’s national team. He was born in <|db_start|> Ko Itakura<|sep|> Birthplace<|db_retrieve|> Yokohama<|db_end|> Yokohama on <|db_start|> Ko Itakura<|sep|> Birth Date<|db_retrieve|> January 27, 1997<|db_end|> 27th January 1997. His father is from <|db_start|> Ko Itakura<|sep|> Father’s Origin<|db_retrieve|> Japanese<|db_end|> Japanese and his mother is from <|db_start|> Ko Itakura<|sep|> Mother’s Origin<|db_retrieve|> Japanese<|db_end|> Japanese. When he was young, he played baseball but after watching an exhibition match against a professional baseball club, he decided to become a footballer. In <|db_start|> Ko Itakura<|sep|> Joined Club Year<|db_retrieve|> Kawasaki Frontale<|db_end|> 2013, he joined J1 League side Kawasaki Frontale. However, he could not play many matches behind Shusaku Nishikawa until <|db_start|> Ko Itakura<|sep|> First Match as Starter<|db_retrieve|> Uruguay<|db_end|> September when he debuted at right back against Uruguay. After that, he became a regular player under manager <|db_start|> Ko Itakura<|sep|> Manager Under Whom Became Regular Player<|db_retrieve|> Japan national team<|db_end|> Shinji Ono. On <|db_start|> Ko Itakura<|sep|> Debut Date<|db_retrieve|> June 17, 2019<|db_end|> 17 June 2019, he debuted with Japan national team against Chile during the <|db_start|> Ko Itakura<|sep|> Competition Debut<|db_retrieve|> Uruguay<|db_end|> 2019 Copa América. Career statistics. “Updated to end of 2018 season”. National team career. In August 2016, Itakura was elected to the <|db_start|> Ko Itakura<|sep|> U-23 Selection<|db_retrieve|> Japan U-20 national team<|db_end|> Japan U-20 national team for the <|db_start|> Japan U-20 national team<|sep|> Tournament Participation<|db_retrieve|> 2017 U-20 World Cup<|db_end|> 2017 U-20 World Cup. At this tournament, he played all 4 matches as left back of three back defense. In May 2019, he was selected for the <|db_start|> Ko Itakura<|sep|> Senior Squad Selection<|db_retrieve|>